OpenStudy (anonymous):

If L is the line having x-intercept of -1 and y-intercept of 3, complete the equation of L. I know that the y-intercept takes the place of b in y=mx +b. So, I have y=mx+3 so far. I'm just not sure where to substitute in my x-intercept.

OpenStudy (anonymous):

Use Slope inter cept form

OpenStudy (anonymous):

x/a + y/b = 1

OpenStudy (anonymous):

x/-1 + y/3 = 1

OpenStudy (anonymous):

-3x + y = 3

OpenStudy (anonymous):

Oh okay, I gotcha. Sorry it took me so long to reply, I was working it out :) Thanks!

OpenStudy (anonymous):

Welcome..)

OpenStudy (anonymous):

Wait. Actually. I think I did something wrong, I got y=3x-3

OpenStudy (anonymous):

x/-1 = -x

OpenStudy (anonymous):

Well this is what I did.. -3(x/-1 +y/-3)=1*-3 3x+y=-3 y=-3x-3 Oh crud. Well now I have two different answers..

OpenStudy (anonymous):

Oh wait, I just forgot the (-) sign the first time I wrote it, so I still got my same answer. It's just different from yours

OpenStudy (anonymous):

lol...+y/-3 hmmm.. it shuld be y/3

OpenStudy (anonymous):

Oh geeze... Whoops. Thanks :) Silly me!